OpenStudy (anonymous):

write the equation of the line that passes through (-5,8) and (-3,-8)

OpenStudy (anonymous):

super simple give me a sec to find the formula

OpenStudy (anonymous):

Oh man I accidently refreshed the page with all the explanations.

OpenStudy (anonymous):

accidentally*

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

I just did this in class ;o

OpenStudy (anonymous):

Alright use the formula \[m= \left( y2-y1/x2-x1 \right)\] thats the formula to find slope. m= your slope. so now use the slope-intercept form and plug in one of the points: \[y-y1= m(x-x1)\]

OpenStudy (anonymous):

1) Find the slope of that line using m = (y2-y1)/(x2-x1) Then plug in the slope (m) into the formula y-y1=m(x-x1) y and x are not variables for you to sub in numbers, it is just variables.. letters. y1 and x1 are the variables where you sub in your numbers from one of the points, pick which ever coordinate you like. Depending on what your textbook or teacher wants, put it in the correct form..

OpenStudy (anonymous):

Alright so you're give two points and you know to find the slop is the change in y over the change in x. So [(8) - (-8)]/[(-5)-(-3)] = 16/ -2 = -8 So now that we know that our slope is m=-8, remember that y=mx+b equation? That equation will come in handy cause that's how you'll find the equation of a line. Let's just put in what we have so far which is: y = (-8)x +b Since two points are given and it says that the line passes through both points, pick one so you could have the values for the x and y values. (We'll choose (-3,-8)) After plugging it in, we need to solve for b (the point where the line passes through the y-axis) y = -8x +b (-8) = (-8)(-3) + b (-8) = 24 + b b = -32 Therefore your line would be: y = -8x -32

OpenStudy (anonymous):

slope*